OpenStudy (anonymous):

What does (1/2x^-1/2)-x^1/2 simplify to?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Is it: \[\left( \frac{ 1 }{ 2x^{-1/2} } \right)-x^{1/2}\] Or: \[\left( \frac{ 1 }{ 2 }x^{-1/2} \right)-x^{1/2}\]

OpenStudy (anonymous):

In any case, apply these rules: \[a^{-b} = \frac{ 1 }{ a^b }\] \[x^{1/2} = \sqrt{x}\]

OpenStudy (anonymous):

\[\frac{1}{2 \sqrt{x}}-\sqrt{x}=\frac{1-2 x}{2 \sqrt{x}} \]